### v2

@@ 3.5 Scalability
4 FedGraph Usability
In this section, we first an example of how to use the library, then show the configuration of homomorphic encryption and measurement dashboard.
4.1 FedGraph Use Example
Researchers and applied scientists can easily install the library, set up configurations, and then start training.
{python}
## Installation pip install fedgraph
After installing the library, the user can run the experiment with the following steps:
Task specification: The user declares the task on which the runner will focus. As we have involved multiple graph learning algorithms that cannot share the same configuration, we require the user to further specify the algorithm they want to deploy instead of manually modifying the configuration file.
Configuration loading: The configuration file contains all the necessary information needed for the training, including dataset details, model hyperparameters, and some general training options. The configuration is set up for different tasks and algorithms.
Data loading: Once started running, a build-in module data_loader is used for loading the data if a correct set of configurations is supplied. We also provide the data interface to support customized datasets.
Distributed graph training: Once the data gets loaded into different clients, the training experiment starts by calling the general runner run_fedgraph.
The implementation details of the API usage are demonstrated in the following code block.
{python}
from fedgraph import data_process from fedgraph import federated_methods import attridict
config = ”dataset”: ”cora”, ”fedgraph_task”: ”NC”, ”method”: ”FedGCN”, ”global_rounds”: 100, ”local_step”: 3, ”learning_rate”: 0.5, ”n_trainer”: 2, ”num_layers”: 2, ”num_hops”: 2, ”iid_beta”: 10000, ”gpu”: False, ”encrypted_computation”: True, ”logdir”: ”./runs”,
config = attridict(config) data = data_process.data_loader(config) federated_methods.run_fedgraph(config,data)
4.2 Homomorphic Encryption Configuration
4.2.1 Parameter Configuration
Parameter Default Value Description Range
scheme CKKS Encryption scheme type N/A
security level 128 Bit security level 128, 192, 256
FedGCN (plaintext) N/A N/A N/A Cora 13.29 59.21 0.783 $\pm$ 0.07
Table 3: Microbenchmark of FedGCN under Homomorphic Encryption with different CKKS scheme parameters.
4.2.2 Microbenchmark
We then provide the microbenchmark of HE on federated graph training in Table 3. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings in Cora. For CKKS parameters, we evaluate different polynomial modulus degrees (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for the encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds. The complete version can be found in Table 6 in Appendix.
4.3 System Measurement Dashboard
Figure 8: FedGraph Monitoring System.
5 System Benchmark Results
In this section, we provide the system benchmark for graph classification, node classification, link prediction, and system scalability.
@@ 5 System Benchmark Results
Figure 9: Accuracy, Communication Cost and Train time of Different Graph Classification Algorithms (10 clients)
5.1 Graph Classification
@@
The graph classification experiment aims to evaluate the effectiveness of various federated learning algorithms in classifying graph data across diverse datasets, including IMDB-BINARY, IMDB-MULTI, MUTAG, BZR, and COX2. Each dataset represents different graph structures, such as molecular and social networks, providing a comprehensive testbed for algorithm performance.
In Figure 9, our evaluation process spans 200 rounds, accuracy results indicate that GCFL+ and GCFL+dWs consistently achieve higher accuracy across datasets, with MUTAG and COX2 showing the best results across all algorithms. FedAvg and GCFL exhibit moderate performance, with accuracy varying based on the dataset’s complexity. GCFL+ and GCFL+dWs have relatively high training times, particularly on complex datasets like MUTAG and BZR. FedAvg shows the shortest training time across datasets, making it more suitable for time-constrained applications.
As shown in the right panel, GCFL incurs the highest communication cost, particularly for BZR, which highlights a trade-off between communication efficiency and algorithmic complexity. This suggests that while GCFL may offer performance advantages, it does so at the expense of higher data exchange. In contrast, FedAvg demonstrates the lowest communication cost, making it a more suitable choice for scenarios with network constraints where minimizing communication is essential.
5.2 Node Classification
In our node classification experiments, we evaluate the impact of different label distribution metrics using the IID Beta setting. We utilize varying values of $\beta$ (10, 100, 10000) to control the data distribution, where larger $\beta$ values represent more IID (independent and identically distributed) scenarios, and smaller values indicate increasingly non-IID distributions. This approach allows us to simulate diverse federated environments and study the algorithms’ performance under varying data conditions.
Our experiments span multiple datasets like Cora, Citeseer, ogbn-arxiv. With ogbn-products and ogbn-papers100M serving as large-scale benchmarks, helping to further understand model scalability.
We also assess several node classification algorithms, including FedAvg and FedGCN, across different batch sizes. This analysis provides insight into the algorithms’ adaptability and efficiency with both IID and non-IID data.
In terms of system monitoring, we tracked key performance metrics across 10 training nodes and a server node. These metrics include test accuracy, training time, memory consumption, and network usage across both pretrain and train phases.
Figure 11: Training Time, Communication Cost, Test Accuracy on Ogbn-Arxiv in Large Number of Clients
Figure 12: Training Time, Communication Cost, Test Accuracy on Ogbn-Papers100M in Different Batch Sizes (800 rounds)
Figure 13: Comparison of FedGCN Execution Time (top) and Communication Cost (bottom) with HE across Different Datasets
@@ 5.2 Node Classification
Figure 14: Training time, Communication Cost, AUC Comparison of Different LP Algorithms
5.3 Mini-Batch for Large Dataset
5.4 Scalability: Increase Number of Clients
To better test scalability and fit real-world data, we increase the number of clients to 1000. In Figure 11, we observe that as the number of clients increases, the overall training time grows significantly due to added communication overhead and the need for increased synchronization among clients.
As we scale from 10 to 1000 clients under a fixed IID Beta value, there is a small decline in accuracy, likely due to the increased data heterogeneity each client possesses. The communication cost also escalates notably with more clients, highlighting the trade-off between parallelism and efficiency in federated settings. This experiment underscores the system’s ability to handle large-scale client distributions while revealing the resources required to maintain accuracy and efficiency.
5.5 Scalability: Real-World Dataset with the 200 countries population
In our experiments on the ogbn-papers100M dataset over 50GB, we used Hugging Face to store and split the dataset, utilizing the Trainer class to manage local loading for each client. We distributed the data across 195 clients, adjusting the node count for each client to fit a power-law distribution based on country population sizes. This approach mimics real-world federated environments, where larger clients receive more data, accurately reflecting varying data availability. Figure 12 illustrates three key metrics as we vary the batch size (16, 32, and 64).
While training times are consistent across batch sizes, pretrain times increase proportionally with batch size, indicating a need for efficient preprocessing management.
Memory usage remains stable (about 17.5GB), showing the model’s memory footprint is well-optimized. Test accuracy is largely unaffected by batch size, with a slight increase from 16 to 32 that levels off at 64, suggesting flexibility in batch size selection without compromising accuracy.
5.6 Homomorphic Encryption
We conduct microbenchmarks through different selections of parameters and methods to evaluate the performance metrics and trade-offs of HE on FedGCN in Table 3. When comparing among datasets, we observe that HE maintains equivalent accuracy across different parameter selections, as long as they satisfy the modulus requirement. If a smaller than required parameter size is used, the accuracy drops sharply which indicates not valid encryption.
We present Figure 13 to visualize the overhead that HE adds to FedGCN. It illustrates the execution time comparison between plain-text and encrypted versions across different datasets. The computational overhead increases particularly with larger datasets. When comparing FedGCN running across different datasets, feature aggregation takes up a larger portion of communication cost with increasing dataset size. This could be due to the difference in feature and model parameter aggregation in FedGCN, as larger datasets usually have more complex feature representations that require more extensive encryption overhead.
5.7 Link Prediction
In the federated link prediction setting, each client holds region-specific data, respecting privacy constraints by avoiding raw data sharing. The algorithms are assessed on AUC for accuracy, training time for computational efficiency, and communication cost for network efficiency in a distributed, region-specific setup. In Figure 14, the experiment shows that STFL and FedLink achieve the highest AUC values across datasets, with StaticGNN and 4D-FED-GNN+ performing moderately well, but notably lower for the simplest dataset (US). For train time, FedLink and STFL exhibit the longest train times, especially on the largest dataset. 4D-FED-GNN+ has the shortest training times, showing efficiency for quicker iterations. FedLink has the highest communication cost, particularly on the largest dataset, while StaticGNN+ maintains the lowest cost across all datasets, making it more network-efficient.
@@ 6 Conclusion
In this paper, we introduce FedGraph, a Python library designed for benchmarking federated graph learning algorithms. Unlike prior evaluation platforms, FedGraph is designed to specifically support multiple FGL algorithms, facilitating their comparisons under different system, algorithm, and dataset configurations. FedGraph supports truly distributed training and homomorphic encryption, allowing for real-world evaluation of communication and computing costs for private FGL. The FedGraph API and system profiler facilitate easy use by machine learning researchers, and the library also allows users to easily introduce their own datasets and algorithms for evaluation. We showcase FedGraph’s ability to support FGL algorithm comparisons and private FGL on large graphs with 100 million nodes, demonstrating that it enables real-world FGL evaluation.
@@ Appendix A FedGraph Code Structure
Appendix B Run_fedgraph
@@
def run_fedgraph(args, data): if args.fedgraph_task == ”NC”: run_NC(args, data) elif args.fedgraph_task == ”GC”: run_GC(args, data) elif args.fedgraph_task == ”LP”: run_LP(args)
@@ Appendix D Runner Workflow
D.1 Runner Workflow for Graph Classification Task
@@
D.2 Runner Workflow for Link Prediction Task
@@
Appendix E Microbenchmark for FedGCN
We present the full version of Microbenchmark for FedGCN in Table 6.

### v3

@@ 3.4 Optimizations for Scalability
4 Case Study: Communication and Computation Efficient Federated Node Classification with Low Rank Feature Compression
Efficient communication is a critical challenge in federated graph learning, especially when handling large graphs or deploying privacy-preserving methods such as homomorphic encryption. To address this, we implement low-rank feature compression within FedGraph that significantly reduces communication overhead while preserving model accuracy.
4.1 Architecture Support in FedGraph
FedGraph’s modular design enables the seamless integration of low-rank compression. It separates the pre-training feature aggregation and model training phases, allowing different optimization strategies at each stage. In this case study, we apply HE in both pre-training and model training aggregations, while the low rank compression is applied during pre-training. The HE interface, integrated across multiple algorithms, naturally supports low-rank encrypted aggregation due to its additive structure.
4.2 Low Rank Method for Pre-Train Feature Aggregation
In FedGCN, pre-training communication involves aggregating feature information across clients for nodes with cross-client edges, which may be encrypted and securely aggregated in order to preserve privacy (Yao et al., 2024). We implement a low-rank method for pre-train feature aggregation using client-side projection of feature information to reduce communication cost. The server first generates a random projection matrix $\mathbf{P}\in\mathbb{R}^{d\times k}$ where $k\ll d$ is the specified rank (e.g., $k=100$). The server then distributes $\mathbf{P}$ to all clients. Each client $i$ computes the projected feature matrix $\mathbf{\hat{X}}_{i}=\mathbf{X}_{i}\mathbf{P}$ where $\mathbf{\hat{X}}_{i}\in\mathbb{R}^{n_{i}\times k}$ with $n_{i}$ nodes and then sends $\mathbf{\hat{X}}_{i}$ to the server. The server aggregates these low-rank projections $\mathbf{\hat{X}}_{agg}=\sum_{i=1}^{m}\mathbf{\hat{X}}_{i}$, where $m$ is the number of clients, and distributes the result back to clients. To prevent potential reconstruction of another client’s original features, the projection matrix $\mathbf{P}$ can also be encrypted before being distributed. This adds a layer of protection against inversion attacks on the shared aggregated features. This approach significantly reduces communication overhead in both directions due to only communicating low-rank information. Since our HE interface supports addition on encrypted data, it can safely perform the aggregation procedure on encrypted projected features, preserving client privacy.
4.3 Performance Evaluation
We evaluate the impact of low-rank compression on communication cost, training time, and accuracy in FedGCN using the Cora dataset. Compression ranks range from the full 1433-dimensional features to rank 100, yielding up to 93% reduction. As shown in Figure 7, HE alone introduces significant communication overhead, especially in pre-training, but this is substantially mitigated by applying low-rank projection. Across both plaintext and HE settings, accuracy remains stable even as rank decreases, demonstrating the effectiveness of low-rank compression for communication-efficient and privacy-preserving FGL.
Figure 7: Comparison of communication cost (first and second plots) and training time (third and fourth plots) for FedGCN with different low-rank compression settings. We show results with plaintext and with HE. Each bar represents the total cost divided into pre-training (blue) and training (green) phases, with accuracy plotted as an orange line to show performance trade-offs.
5 Benchmarking FedGraph on Graph Learning Tasks and Scalability
In this section, we benchmark FedGraph’s performance on graph classification, node classification, and link prediction tasks. We then benchmark FedGraph’s performance as we scale the size of the dataset, demonstrating the library’s ability to scale well. Appendix G provides additional results.
5.1 Federated Graph Classification
@@
Figure 8: Accuracy (left), Training Time (middle), and Communication Cost (right) Comparison across Federated Graph Classification Algorithms (10 clients).
@@
The graph classification experiment evaluates the performance of various federated learning algorithms on five benchmark datasets: IMDB-BINARY, IMDB-MULTI, MUTAG, BZR, and COX2. These datasets span diverse graph structures, including social and molecular networks, providing a comprehensive testbed for assessing algorithm effectiveness.
As shown in Figure 8, the evaluation spans 200 training rounds across five datasets. GCFL+ and GCFL+dWs consistently achieve the highest accuracy, particularly on BZR and COX2. However, as illustrated in the middle and right plots, these gains come at the cost of significantly higher training time and communication overhead, especially on complex datasets like IMDB, reflecting their greater computational and communication demands. In contrast, FedAvg offers the lowest communication cost and the shortest, most consistent training time across all datasets, making it a practical choice for latency- and bandwidth-constrained environments.
5.2 Federated Node Classification
Figure 9: Accuracy (left), Training Time (middle), and Communication Cost (right) Comparison across Federated Node Classification Algorithms under $\beta=10000$ (IID data distribution).
We evaluate FedAvg and FedGCN on four node classification benchmarks: Cora, Citeseer, PubMed, and Ogbn-Arxiv. As shown in Figure 9 (left), FedGCN consistently achieves higher accuracy than FedAvg across all datasets. However, this improvement comes at the cost of significantly longer training times, particularly on larger datasets such as Ogbn-Arxiv in Figure 9 (middle).
The communication cost breakdown (Figure 9, right) reveals the source of this overhead: FedGCN incurs substantially higher pre-training communication costs compared to FedAvg. Notably, the observed communication costs closely match the theoretical values, validating the efficiency modeling of these algorithms. This is especially pronounced for Ogbn-Arxiv, where the pre-training phase dominates the total communication volume. While FedAvg has relatively balanced communication between pre-training and training phases, FedGCN’s architecture leads to a disproportionate communication burden during the pre-training aggregation of features across clients. This imbalance helps explain the increased total training time observed for FedGCN on large-scale datasets.
5.3 Federated Link Prediction
@@
Figure 10: AUC (left), Training Time (middle), and Communication Cost (right) Comparison across Federated Link Prediction Algorithms (10 clients).
In the federated link prediction setting, we simulate a scenario where each client holds region-specific data from the Foursquare Global-scale Check-in Dataset (Yang et al., 2016), covering three geographic configurations: US; US and BR; and US, BR, ID, TR, and JP. This setup respects data privacy by avoiding raw data sharing across regions. We evaluate four algorithms, 4D-FED-GNN+, FedLink, STFL, and StaticGNN, across three key metrics: AUC for predictive accuracy, training time for computational efficiency, and communication cost for network efficiency.
As shown in Figure 10, FedLink and STFL achieve the highest AUC scores across all datasets, while StaticGNN and 4D-FED-GNN+ perform moderately well, but show lower AUC on the simpler US-only dataset. In terms of training time, FedLink and STFL incur the highest costs, particularly on the largest regional dataset. In contrast, 4D-FED-GNN+ demonstrates the shortest training times, suggesting suitability for fast iterative training. For communication efficiency, FedLink incurs the highest overhead, especially on large datasets, while StaticGNN consistently exhibits the lowest communication cost, making it the most network-efficient among the evaluated methods.
Figure 11: Accuracy Across Global Training Rounds and CPU Usage for Node Classification Tasks.
5.4 System Performance Monitoring and Resource Utilization
We monitor system performance across 10 training nodes and a central server node using Grafana, which visualizes real-time metrics collected via Prometheus. Figure 11 provides an overview of system behavior during node classification experiments. The first three plots show global test accuracy across training rounds for the Cora, Citeseer, and Pubmed datasets. FedGCN demonstrates significantly faster convergence and higher final accuracy than FedAvg across all datasets. Figure 11 (right) shows CPU usage over time, captured via Grafana. Lighter workloads, such as Cora and Citeseer, result in lower and less frequent CPU utilization spikes. In contrast, Pubmed and Ogbn-Arxiv induce higher and more sustained CPU usage due to their larger graph sizes and increased communication overhead. The pattern of usage spikes aligns with scheduled training rounds, reflecting real-time computational demands at scale.
5.5 Real-World Dataset with Realistic Client Data Distribution
Figure 12: Training Time and Test Accuracy on Ogbn-Papers100M under Different Batch Sizes (800 rounds).
We finally evaluate FedGraph at scale on the Ogbn-Papers100M dataset (over 50GB), one of the largest publicly available graph benchmarks. We use Hugging Face for dataset storage and partitioning, with the Trainer class managing local data loading for each client. The data is distributed across 195 clients  [^2] https://huggingface.co/FedGraph, with node counts assigned following a power law distribution based on country population sizes—mimicking realistic federated environments where larger clients hold more data.
As shown in Figure 12, we assess the effect of varying batch sizes (16, 32, and 64) on training time and test accuracy over 800 training rounds. Training time increases modestly with batch size due to additional per-round computation. Test accuracy remains nearly unchanged, with a slight gain from batch size 16 to 32 and a plateau at 64, suggesting flexibility in batch size selection without compromising model performance. Memory usage remains stable at approximately 17.5 GB, demonstrating FedGraph’s scalability and efficient resource handling in large-scale settings.
@@ 6 Conclusion
In this paper, we presented FedGraph, a Python library designed for benchmarking federated graph learning algorithms. Unlike general federated learning platforms, FedGraph supports a diverse set of algorithms and enables systematic comparisons across algorithms, datasets, and system configurations. It features fully distributed training, homomorphic encryption for privacy-preserving scenarios, and a built-in system profiler to measure communication and computation overhead. The modular API allows easy integration of custom datasets and algorithms. Through extensive experiments, including low-rank compression and large-scale training on graphs with up to 100 million nodes, we demonstrate that FedGraph is a practical and scalable tool for real-world FGL evaluation.
While this work focuses on enabling privacy-preserving federated graph learning, future efforts are needed to explore more robust privacy risk assessments, additional optimization strategies, and the inclusion of a broader range of FGL algorithms to expand benchmark coverage and better support industrial deployment.
@@ Appendix A FedGraph Code Structure
Appendix B FedGraph API and Runners
Figure 13: High-level Demonstration of FedGraph API Design
@@ Appendix B FedGraph API and Runners
def run_fedgraph(args, data): if args.fedgraph_task == "NC": run_NC(args, data) elif args.fedgraph_task == "GC": run_GC(args, data) elif args.fedgraph_task == "LP": run_LP(args)
@@ Appendix D Runner Workflow
D.1 Runner Workflow for Node Classification Task
For the node classification task, the user specifies the dataset name in the configuration file. The dataset is preprocessed and partitioned across clients based on the chosen federated setting. The data is accessed directly from local storage or via an API, and then a time window may be generated to support temporal learning tasks.
Dataset extraction and preprocessing are managed by the function dataloader_NC in the module data_process.py, which handles both temporal and static graph datasets. Each client holds a subgraph or node features, and training proceeds in a federated manner.
[…]
D.2 Runner Workflow for Graph Classification Task
@@
D.3 Runner Workflow for Link Prediction Task
@@
Appendix E System Measurement Dashboard
Figure 15: FedGraph Monitoring System.
Appendix F Benchmark Homomorphic Encryption
F.1 Parameter Configuration
Parameter Default Value Description Range
scheme CKKS Encryption scheme type N/A
security level 128 Bit security level 128, 192, 256
F.2 Microbenchmark
We then provide the microbenchmark of HE on federated graph training in Table 5. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings in Cora. For CKKS parameters, we evaluate different polynomial modulus degrees (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for the encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds.
When comparing among datasets, we observe that HE maintains equivalent accuracy across different parameter selections, as long as they satisfy the modulus requirement. If a smaller than required parameter size is used, the accuracy drops sharply which indicates not valid encryption.
@@ F.2 Microbenchmark
Appendix G Additional Experiments
G.1 Mini-Batch for Large Dataset
G.2 Increasing the Number of Clients
Figure 17: Training Time, Communication Cost, Test Accuracy on Ogbn-Arxiv in Large Number of Clients. All experiments run on 10 AWS instances. 1000 trainers take long since it runs sequentially on 10 instances.
To better test scalability and fit real-world data, we increase the number of clients to 1000. In Figure 17, we observe that as the number of clients increases, the overall training time grows significantly due to sequential running on 10 instances, added communication overhead and the need for increased synchronization among clients.
As we scale from 10 to 1000 clients under a fixed IID Beta value, there is a small decline in accuracy, likely due to the increased data heterogeneity each client possesses. The communication cost also escalates notably with more clients, highlighting the trade-off between parallelism and efficiency in federated settings. This experiment shows the system’s ability to handle large-scale client distributions while revealing the resources required to maintain accuracy and efficiency.